the length of a rectangle is 6 more inches than the width. the area of the rectangle is 216 square inches. what are the dimensions of the rectangle





Answer by stanbon(75887)

The moving sidewalk at Chicago's O'Hare International Airport moves at a speed of 2.0 feet per second. Walking on the moving sidewalk, Nancy Killian walks 120 feet in the same time that it takes her to walk 52 feet without the moving sidewalk. How fast does Nancy walk?
----------
Let Nancy's walking rate be "n".
----------------------------------------
With moving sidewalk DATA;
distance = 120 ft ; rate = (n+2) ft/sec ; time = 120/(n+2) seconds
-----------------------------------
Without moving sidewald DATA:
distance = 52 ft ; rate = n ft/sec ; time = 52/n seconds
-----
Equation:
time = time
120/(n+2) = 52/n
----
120n = 52n+104
68n = 104
n = 104/68 = 1.53 ft.sec (Nancy's walking rate)
===============================================
Cheers,
Stan H.